# Gaze-LLE: Gaze Target Estimation via Large-Scale Learned Encoders

### v1

@@ Abstract
We address the problem of gaze target estimation, which aims to predict where a person is looking in a scene. Predicting a person’s gaze target requires reasoning both about the person’s appearance and the contents of the scene. Prior works have developed increasingly complex, hand-crafted pipelines for gaze target estimation that carefully fuse features from separate scene encoders, head encoders, and auxiliary models for signals like depth and pose. Motivated by the success of general-purpose feature extractors on a variety of visual tasks, we propose Gaze-LLE, a novel transformer framework that streamlines gaze target estimation by leveraging features from a frozen DINOv2 encoder. We extract a single feature representation for the scene, and apply a person-specific positional prompt to decode gaze with a lightweight module. We demonstrate state-of-the-art performance across several gaze benchmarks and provide extensive analysis to validate our design choices. Our code is available at: http://github.com/fkryan/gazelle.
*Equal advising.
@@ 1 Introduction
Gupta et al.[22] ${}_{\text{IMAGE}}$ Original 0.933 0.134 0.071
Trained DINOv2 ViT-B 0.912 0.155 0.090
@@
Table 1: Existing gaze architectures do not leverage features from large transformer models effectively. We replace the scene encoder in 3 existing open source methods with the DINOv2 ViT-B backbone and evaluate on GazeFollow (see Appendix for details). Using DINOv2 does not improve performance—whether or not its parameters are frozen.
@@ 2 Related work
The dominant approach to gaze target estimation is a multi-branch fusion approach, in which an initial encoder is followed two or more analysis branches that work in parallel to extract specific cues for gaze estimation. These branches converge in a fusion module which produces an integrated representation which is then decoded into the output heatmap, with end-to-end training of the entire pipeline. In contrast, the goal of this paper is to demonstrate that multi-branch architectures are unnecessary and inefficient. We show that SotA performance can be obtained by processing the feature representation produced by a frozen foundational visual encoder, using a novel decoder architecture which incorporates the unique properties of gaze estimation. We are the first to demonstrate the feasibility of using a frozen large-scale encoder for this task, and the design of our decoder architecture is novel relative to prior gaze estimation works.
@@
Among the multi-branch architectures, the two-branch approach of Tafasca et. al. [59] is the most closely-related to this paper. They use a large transformer-based backbone for the scene analysis branch, which receives the head branch representation as input. While they initialize their backbone with pretrained weights, it is still trained end-to-end, with the head branch producing a specialized feature representation. In contrast, we demonstrate that 1) head analysis can also be directly integrated into the decoder, eliminating the need for a head branch and further simplifying and streamlining the architecture, and 2) frozen large-scale foundational encoders give superior performance with two orders of magnitude fewer learned parameters. Gaze-LLE produces higher accuracy on all datasets (see Table 3).
Some prior works [64, 65, 63] have explored an alternative formulation of gaze target estimation as a set detection problem, where a model based on DETR [3] jointly predicts the location of all heads and their accompanying gaze targets. This avoids the need for a separate head detection step, but creates additional complications and existing works require the input of ground truth gaze data at testing time - prevnting comparison with most methods (see Appendix). Other prior works have estimated 3D gaze direction from facial appearance [31, 70, 15] without identifying the gaze target, and some early approaches estimated head orientation [56, 69] to identify gaze targets.
Another area of gaze behavior recognition involves the joint analysis of multi-person social gaze behaviors, such as shared attention (when 2 people are looking at the same gaze target) [24, 14, 57, 44, 47, 54], mutual gaze (when a pair of people is looking at each other) [38, 39, 46, 40, 11, 41], and other gaze-related social structures [15, 17]. Such analysis can be used in assessing and understanding social behaviors for conditions like autism [35, 9, 7], and may benefit from a gaze estimation approach like ours where social context is naturally encoded in a shared scene representation.
@@ 3.2 Key Design Decisions for Foundation Models
Our first finding is that a straightforward substitution of DINOv2 into prior gaze architectures leads to consistently poor performance. In Tab. 1, we show the result of swapping the scene encoder in three open source gaze estimation methods with DINOv2 (see Appendix for more results). Whether or not we finetune the DINOv2 backbone, it is outperformed by the supposedly “weaker” backbones in these prior works. This is not necessarily surprising, as prior gaze works [64, 65, 30] have found ResNet-50 to sometimes outperform more powerful architectures. But this finding reinforces the need for our Gaze-LLE solution. To gain further insight, we conducted a set of experiments on GazeFollow with a simple baseline: extract scene and head features with a frozen DINOv2, concatenate the results, and decode into a gaze heatmap (see Appendix for full details). With this baseline, we quantify the impact of our three key architectural choices: integration of the head position, design of feature decoding, and use of a head branch (Tab. 2).
@@
Prior works use a separate encoder that inputs a crop of the head, which is useful for understanding gaze direction. We hypothesize that a large-scale encoder like DINOv2 already captures gaze direction in its representation. We compare performance with and without a head branch and find it to be nearly the same when using a transformer-based decoder (d v.s. f). Notably, this doesn’t occur with the conv decoder (c v.s. e), indicating that the relevant features are already there, but we need a transformer’s global information propagation to extract them. This experiment motivated our novel head prompting design in Gaze-LLE.
@@
Table 3: Gaze target estimation results on GazeFollow and VideoAttentionTarget. We report the number of learnable parameters for each model, and if auxiliary models are used for inputs: I is image, D is depth, and P is pose, O is objects, and E is eyes. (Parameter estimate. Our reimplementation, see Appendix. Metric re-evaluated to match benchmark’s calculation protocol [9].)
@@ 4.1 Main Results
Tab. 3 compares the performance of Gaze-LLE with existing methods on GazeFollow and VideoAttentionTarget. Our model achieves SotA on the AUC and L2 metrics for both datasets, while using only using a single image encoding branch and with a small fraction of the learnable parameters of prior approaches. We include results for both DINOv2 ViT-B and DINOv2 ViT-L, observing that the ViT-L backbone produces stronger results while ViT-B is still sufficient for obtaining SotA. Importantly, our ViT-B model outperforms Tafasca et al. [59], which also uses a ViT-B backbone, but trains the backbone end-to-end along with a separate head branch and large DPT decoder. Our approach’s design effectively leverages the power of the pretrained backbone, achieving stronger results with $\approx 2\%$ of the learned parameters. An additional benefit of our approach is dramatically reduced training time. As illustrated in Fig. 4, our model converges much faster than prior methods, achieving SotA in less than 1.5 hours on a single RTX4090.
On VideoAttentionTarget’s in/out of frame prediction task, we obtain second-best results; however our method obtains the best results when considering all metrics together. Our approach’s strong, SotA results validate our hypothesis that DINOv2 features do indeed capture appropriate information for gaze target estimation, and that a single-stream design can outperform multi-stream architectures with additional modalities. To assess our model’s performance on a more specialized dataset, we include results on ChildPlay in Tab. 4. Our ViT-B and ViT-L models both achieve SotA results across all AUC, L2, and AP. We also include the benchmark’s P.Head metric [58], which assesses the precision predicting when gaze targets lie within a head bounding box; however due to its reliance on detections which aren’t always accurate, we find this metric tends not to correlate with the others.
@@ 4.2 Analysis
While we use DINOv2 in our main experiments, Gaze-LLE can be used with any backbone. Table 6 reports our model’s performance with different pretrained encoders. The supervised [55] and MAE [23] models are pretrained on ImageNet-1k [10], while CLIP [48] and DINOv2 [45] are trained on much larger data sources. Unsurprisingly, DINOv2, which is the state-of-the-art for general-purpose feature extraction on dense downstream tasks, performs best, but CLIP also achieves strong results. As new backbones are developed, Gaze-LLE provides a framework for adapting them to gaze estimation.
@@
We also consider integrating the head position as its own token during attention as an alternative to our added position embedding $p_{\text{head}}$, inspired by works in point tracking and segmentation that represent positional queries as tokens [29, 32]. We construct a head position token, $t_{\text{pos}}$, by sampling the position embedding $P$ at the head bounding box’s center point and summing this with a learned embedding. We concatenate this token to the scene token list, $S$. To fuse positional information with the scene features, we consider two types of attention in the transformer layers: self attention across the full token list, which updates both $S$ and $t_{\text{pos}}$, and cross attention from the scene tokens to the position token, which updates only $t_{\text{pos}}$. Finally, we consider 3 methods for producing the heatmap from $S$ and $t_{\text{pos}}$: a 2-layer convolutional decoder on $S$ (as used in in our default method), replacing the second convolutional layer with the dot product between $t_{\text{pos}}$ and the scene feature map (like transformer segmentation methods [6, 5, 32]), and directly regressing the $64\times 64$ heatmap with a 2-layer MLP. We show results in Tab. 7 and find that given the right settings, this position token can be made almost as effective as our default embedding approach—however, with the added benefit of potentially being able to decode multiple head positions at the same time (with each new position being another token). For our lightweight decoder, additional head locations already add negligible compute (see Appendix), but future work with heavier gaze decoders may benefit from this token design.
@@
We also investigate where in the decoding process to inject the head prompt. To explore the tradeoff between performance and the reduction of person-specific computation, we move the head prompting to later transformer layers in our gaze decoder, and show that this yields only small performance drops (see Table 8(b)). This demonstrates SotA results on GazeFollow with the only per-person computation taking place in the final transformer layer and 2 convolutional layers. The ability to minimize person-specific computation while maintaining strong performance could provide efficient scaling for multi-person gaze analysis. We also investigate the performance when head detections from YOLOv5 are used instead of ground truth head bounding boxes, and observe almost no degradation. See the Appendix for the details.
@@ 5 Discussion
By leveraging a frozen encoder without end-to-end training, our performance is inherently tied to the encoder quality. We find it is important to select an encoder trained on a large, diverse dataset with a dense objective (see Tab. 6). Additionally, while our method is reasonably efficient ($>$50 fps on an RTX4090, see Appendix), the overall efficiency depends on the use of a large encoder, which may pose a challenge for embedded systems. We note that recent approaches that depend on auxiliary transformer-based depth/pose models also experience this limitation. However, as stronger, faster general-purpose feature extractors become available, Gaze-LLE provides a way to harness them for gaze estimation.
@@ Acknowledgments
The authors thank Stefan Stojanov for helpful discussions, and the members of the Hoffman Lab and Rehg Lab for their feedback. This work is in part supported by NIH R01 MH114999, NSF #2144194, and the NSF Graduate Research Fellowship under Grant No. DGE-2039655. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
@@
  • 10 Performance with Estimated Head Bounding Boxes
  • 11 Reimplementation of Horanyi et al.
  • 12 Additional Ablation Studies
  • 13 LoRA Backbones
  • 14 Additional Visualizations & Failure Modes
@@ 6 Integration of DINOv2 into Existing Methods
In this section, we provide further details on our experiments in Table 1, which integrate DINOv2 into three existing methods: Chong et al. [9], Miao et al. [42], and Gupta et al. [22].
@@
In this section, we provide further details about our experiments in Section 3.2 that investigate early vs. late head position integration, transformer vs. convolutional decoding, and head & scene branch vs. scene-branch only design.
@@ 8 Comparison to Detection Methods
A set of recent works formulate gaze target estimation as a set detection problem, jointly predicting a set of head bounding boxes and their corresponding gaze locations [64, 63, 65]. We exclude these works from our main comparisons in Section 4.2 due to differences in the evaluation setting, as these methods perform bipartite matching using the ground truth gaze targets at test time. In this section, we provide further details about the difference in evaluation setting, and provide quantitative comparison with Tonini et al. [63] in our setting using their open source codebase.
@@
At inference time, these methods use the same matching cost $\mathcal{L}_{\text{cost}}$ to determine which predicted instances are evaluated against which ground truth instances, and use this to calculate the gaze performance metrics (e.g. heatmap AUC, L2 distance). This is inherently a different evaluation setting than ours because the ground truth gaze labels are used at inference time to retrieve the predicted instances that are compared against the ground truth instances. Because the fixed number of predicted instances, $N=20$, is much higher than the typical number of ground truth instances per image, a model can predict multiple instances with the same head bounding box, but different gaze targets (see Fig. 6 for visual examples of this). In this case, the matching algorithm will match each ground truth instance to the predicted instance with the closest gaze and calculate the gaze metrics between these pairs, discarding the extra incorrect predictions from evaluation. In this way, the gaze performance metrics alone do not penalize overdetection. They characterize recall by assessing the accuracy of the predicted instances that are closest to the ground truth, but do not assess precision by penalizing the model for predicting additional instances with incorrect gaze targets. Additionally, the matching algorithm does not enforce that the ground truth instance is matched to a detection with a similar predicted head bounding box; if the heatmap loss dominates the matching cost, an instance may be selected based only on similarity between the predicted gaze heatmap and ground truth (see Fig 6 for examples). Thus, the model does not need to correctly associate people with their respective gaze targets to achieve high performance. For these reasons, the gaze metrics in this evaluation setting are not a direct comparison against our work and prior methods that follow the traditional problem formulation.
@@
We show the reported results of the 3 detection-based methods and our results on GazeFollow in Table 14. To quantitatively characterize the difference in evaluation setting, we also re-evaluate Tonini et al.’s [63] method on GazeFollow with the ground truth gaze label removed from the matching cost, using their published codebase. We alter the matching cost from Equation 6 to exclude the ground truth gaze label. This altered matching cost is defined as:
@@
We visualize the output instances of Tonini et al.’s default matching algorithm that uses ground truth gaze as part of the cost function, and our altered matching algorithm that does not use ground truth gaze in Figure 6. The first two rows demonstrate cases where the matching algorithm chooses an instance with a predicted bounding box that is not associated with the correct person; the heatmap loss dominates the matching cost. With our altered matching function, an instance with a predicted bounding box for the correct person but incorrect gaze heatmap is retrieved. The third row shows an example of overdetection, where multiple instances are predicted with a head bounding box for the correct person, but different gaze targets. With the use of ground truth gaze during matching, the instance with the most correct heatmap is selected. However, without this ground truth information, the model does not select the best instance and produces an incorrect gaze prediction. These examples visually illustrate the difference in evaluation setting: when ground truth gaze information is used at test time, a model can achieve high performance by producing instances that capture different potential gaze targets and relying on the matching algorithm to retrieve the best instances to evaluate with. However, the gaze metrics do not characterize the model’s ability to determine which of these instances are indeed gaze targets and associate them with the correct person.

### v2

@@ Abstract
We address the problem of gaze target estimation, which aims to predict where a person is looking in a scene. Predicting a person’s gaze target requires reasoning both about the person’s appearance and the contents of the scene. Prior works have developed increasingly complex, hand-crafted pipelines for gaze target estimation that carefully fuse features from separate scene encoders, head encoders, and auxiliary models for signals like depth and pose. Motivated by the success of general-purpose feature extractors on a variety of visual tasks, we propose Gaze-LLE, a novel transformer framework that streamlines gaze target estimation by leveraging features from a frozen DINOv2 encoder. We extract a single feature representation for the scene, and apply a person-specific positional prompt to decode gaze with a lightweight module. We demonstrate state-of-the-art performance across several gaze benchmarks and provide extensive analysis to validate our design choices. Our code and models are available at: http://github.com/fkryan/gazelle.
*Equal contribution.
@@ 1 Introduction
Gupta et al. [23] Original (EfficientNet-B1) 0.933 0.134 0.071
${}^{\text{(image-only)}}$ Trained DINOv2 ViT-B 0.912 0.155 0.090
@@
Table 1: Existing gaze architectures do not leverage features from large transformer models effectively. We replace the scene encoder in 3 existing open source methods with the DINOv2 ViT-B backbone and evaluate on GazeFollow (see Supp. Sec. 6 for details). Using DINOv2 does not improve performance—whether or not its parameters are frozen.
@@ 2 Related work
The dominant approach to gaze target estimation is a multi-branch fusion approach, in which an initial encoder is followed two or more analysis branches that work in parallel to extract specific cues for gaze estimation. These branches converge in a fusion module which produces an integrated representation which is then decoded into the output heatmap, with end-to-end training of the entire pipeline. In contrast, the goal of this paper is to show that SotA performance can be obtained by processing the feature representation produced by a frozen foundational visual encoder, using a novel decoder architecture. We are the first to demonstrate the feasibility of using a frozen large-scale encoder for this task, and the design of our decoder architecture is novel relative to prior gaze estimation works.
@@
Among the multi-branch architectures, the two-branch approach of Tafasca et. al. [61] is the most closely-related to this paper. They use a large transformer-based backbone for the scene analysis branch, which receives the head branch representation as input. While they initialize their backbone with pretrained weights, it is still trained end-to-end, with the head branch producing a specialized feature representation. In contrast, we demonstrate that 1) head analysis can also be directly integrated into the decoder, eliminating the need for a head branch and further simplifying and streamlining the architecture, and 2) frozen large-scale foundational encoders give superior performance with two orders of magnitude fewer learned parameters. Gaze-LLE produces higher accuracy on all datasets (see Tab. 3).
Some prior works [66, 67, 65] have explored an alternative formulation of gaze target estimation as a set detection problem, where a model based on DETR [3] jointly predicts the location of all heads and their accompanying gaze targets. While this avoids the need for a separate head detection step, it requires complex training, and these works use the ground truth gaze at inference time for matching - which is not consistent with practical use cases and prevents comparison with most methods (see Supp. Sec. 8). Other prior works have estimated 3D gaze direction from facial appearance [32, 73, 16] without identifying the gaze target, and some early approaches estimated head orientation [58, 71] to identify gaze targets.
Another area of gaze behavior recognition involves the joint analysis of multi-person social gaze behaviors, such as shared attention (when 2 people are looking at the same gaze target) [25, 15, 59, 45, 48, 56], mutual gaze (when a pair of people is looking at each other) [39, 40, 47, 41, 12, 42], and other gaze-related social structures [16, 18, 10]. Such analysis can be used in assessing and understanding social behaviors for conditions like autism [36, 9, 7], and may benefit from an approach like ours where social context is naturally encoded within a shared scene representation.
@@ 3.2 Key Design Decisions for Foundation Models
Our first finding is that a straightforward substitution of DINOv2 into prior gaze architectures leads to consistently poor performance. In Tab. 1, we show the result of swapping the scene encoder in three open source gaze estimation methods with DINOv2 (see Supp. Sec. 6 for more results). Whether or not we finetune the DINOv2 backbone, it is outperformed by the supposedly “weaker” backbones in these prior works. This is not necessarily surprising, as prior gaze works [66, 67, 31] have found ResNet-50 to sometimes outperform more powerful architectures. But this finding reinforces the need for our Gaze-LLE solution. To gain further insight, we conducted a set of experiments on GazeFollow with a simple baseline: extract scene and head features with a frozen DINOv2, concatenate the results, and decode into a gaze heatmap (see Supp. Sec. 7 for full details). With this baseline, we quantify the impact of our three key architectural choices: integration of the head position, design of feature decoding, and use of a head branch (Tab. 2).
@@
Prior works use a separate encoder that inputs a crop of the head, which is useful for understanding gaze direction. We hypothesize that a large-scale encoder like DINOv2 already captures gaze direction in its representation. We compare performance with and without a head branch and find it to be nearly the same when using a transformer-based decoder (Tab. 2: d v.s. f). Notably, this doesn’t occur with the conv decoder (Tab. 2: c v.s. e), indicating that the relevant features are already there, but we need a transformer’s global information propagation to extract them. This experiment motivated our novel head prompting design in Gaze-LLE.
@@
Table 3: Gaze target estimation results on GazeFollow and VideoAttentionTarget. We report the number of learnable parameters for each model, and if auxiliary models are used for inputs: I is image, D is depth, and P is pose, O is objects, and E is eyes. (Parameter estimate. Our reimplementation, see Supp. Sec. 12. Metric re-evaluated to match benchmark’s calculation protocol [9].)
@@ 4.1 Main Results
Tab. 3 compares the performance of Gaze-LLE with existing methods on GazeFollow and VideoAttentionTarget. Our model achieves SotA on the AUC and L2 metrics for both datasets, while using only using a single image encoding branch and with a small fraction of the learnable parameters of prior approaches. We include results for both DINOv2 ViT-B and DINOv2 ViT-L, observing that the ViT-L backbone produces stronger results while ViT-B is still sufficient for obtaining SotA. Importantly, our ViT-B model outperforms Tafasca et al. [61], which also uses a ViT-B backbone, but trains the backbone end-to-end along with a separate head branch and large DPT decoder. Our approach’s design effectively leverages the power of the pretrained backbone, achieving stronger results with $\approx 2\%$ of the learned parameters. An additional benefit of our approach is dramatically reduced training time. As illustrated in Fig. 4, our model converges much faster than prior methods, achieving SotA results in less than 1.5 hours on a single Nvidia RTX4090.
On VideoAttentionTarget’s in/out of frame prediction task, we obtain second-best results; however our method obtains the best results when considering all metrics together. Our approach’s strong, SotA results validate our hypothesis that DINOv2 features do indeed capture appropriate information for gaze target estimation, and that a single-stream design can outperform multi-stream architectures with additional modalities. To assess our model’s performance on a more specialized dataset, we include results on ChildPlay in Tab. 4. Our ViT-B and ViT-L models both achieve SotA results across all AUC, L2, and AP. We also include the benchmark’s P.Head metric [60], which assesses the precision predicting when gaze targets lie within a head bounding box; however due to its reliance on detections which are not always accurate, we find this metric tends not to correlate with the others.
@@ 4.2 Analysis
While we use DINOv2 in our main experiments, Gaze-LLE can be used with any backbone. Tab. 6 reports our model’s performance with different pretrained encoders. The supervised [57] and MAE [24] models are pretrained on ImageNet-1k [11], while CLIP [49] and DINOv2 [46] are trained on much larger data sources. Unsurprisingly, DINOv2, which is the state-of-the-art for general-purpose feature extraction on dense downstream tasks, performs best, but CLIP also achieves strong results. As new backbones are developed, Gaze-LLE provides a framework for adapting them to gaze estimation.
@@
We also consider integrating the head position as its own token during attention as an alternative to our added position embedding $p_{\text{head}}$, inspired by works in point tracking and segmentation that represent positional queries as tokens [30, 33]. We construct a head position token, $t_{\text{pos}}$, by sampling the position embedding $P$ at the head bounding box’s center point and summing this with a learned embedding. We concatenate this token to the scene token list, $S$. To fuse positional information with the scene features, we consider two types of attention in the transformer layers: self attention across the full token list, which updates both $S$ and $t_{\text{pos}}$, and cross attention from the scene tokens to the position token, which updates only $t_{\text{pos}}$. Finally, we consider 3 methods for producing the heatmap from $S$ and $t_{\text{pos}}$: a 2-layer convolutional decoder on $S$ (as used in in our default method), replacing the second convolutional layer with the dot product between $t_{\text{pos}}$ and the scene feature map (like transformer segmentation methods [6, 5, 33]), and directly regressing the $64\times 64$ heatmap with a 2-layer MLP. We show results in Tab. 7 and find that given the right settings, this position token can be made almost as effective as our default embedding approach—however, with the added benefit of potentially being able to decode multiple head positions at the same time (with each new position being another token). For our lightweight decoder, additional head locations already add negligible compute (see Supp. Sec. 9), but future work with heavier gaze decoders may benefit from this token design.
@@
We also investigate where in the decoding process to inject the head prompt. To explore the tradeoff between performance and the reduction of person-specific computation, we move the head prompting to later transformer layers in our gaze decoder, and show that this yields only small performance drops (see Tab. 8(b)). This demonstrates SotA results on GazeFollow with the only per-person computation taking place in the final transformer layer and 2 convolutional layers. The ability to minimize person-specific computation while maintaining strong performance could provide efficient scaling for multi-person gaze analysis. We also investigate the performance when head detections from YOLOv5 are used instead of ground truth head bounding boxes, and observe almost no degradation. See Supp. Sec. 11 for the details.
@@ 5 Discussion
By leveraging a frozen encoder without end-to-end training, our performance is inherently tied to the encoder quality. We find it is important to select an encoder trained on a large, diverse dataset with a dense objective (see Tab. 6). Additionally, while our method is reasonably efficient ($>$50 fps on an Nvidia RTX4090, see Supp. Sec. 9), the overall efficiency depends on the use of a large encoder, which may pose a challenge for embedded systems. We note that recent approaches that depend on auxiliary transformer-based depth/pose models also experience this limitation. However, as stronger, faster general-purpose feature extractors become available, Gaze-LLE provides a way to harness them for gaze estimation.
@@ Acknowledgments
The authors thank Stefan Stojanov for helpful discussions, and the members of the Hoffman Lab and Rehg Lab for their feedback. Portions of this work were supported in part by NIH R01 MH114999, NSF #2144194, and the NSF Graduate Research Fellowship under Grant No. DGE-2039655. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
@@
  • 10 Comparison to ViTGaze
  • 11 Performance with Estimated Head Bounding Boxes
  • 12 Reimplementation of Horanyi et al.
  • 13 Additional Ablation Studies
  • 14 LoRA Backbones
  • 15 Additional Visualizations & Failure Modes
@@ 6 Integration of DINOv2 into Existing Methods
In this section, we provide further details on our experiments in Tab. 1, which integrate DINOv2 into three existing methods: Chong et al. [9], Miao et al. [43], and Gupta et al. [23].
@@
In this section, we provide further details about our experiments in Sec. 3.2 that investigate early vs. late head position integration, transformer vs. convolutional decoding, and head & scene branch vs. scene-branch only design.
@@ 8 Comparison to Detection Methods
A set of recent works formulate gaze target estimation as a set detection problem, jointly predicting a set of head bounding boxes and their corresponding gaze locations [66, 65, 67]. We exclude these works from our main comparisons in Sec.  4.1 due to differences in the evaluation setting, as these methods perform bipartite matching using the ground truth gaze targets at test time. In this section, we provide further details about the difference in evaluation setting, and provide quantitative comparison with Tonini et al. [65] in our setting using their open source codebase.
@@
At inference time, these methods use the same matching cost $\mathcal{L}_{\text{cost}}$ to determine which predicted instances are evaluated against which ground truth instances, and use this to calculate the gaze performance metrics (e.g. heatmap AUC, L2 distance). This is inherently a different evaluation setting than ours because the ground truth gaze labels are used at inference time to retrieve the predicted instances that are compared against the ground truth instances. Because the fixed number of predicted instances ($N=20$ for HGTTR) is much higher than the typical number of ground truth instances per image, a model can predict multiple instances with the same head bounding box, but different gaze targets (see Fig. 6 for visual examples of this). In this case, the matching algorithm will match each ground truth instance to the predicted instance with the closest gaze and calculate the gaze metrics between these pairs, discarding the extra incorrect predictions from evaluation. In this way, the gaze performance metrics alone do not penalize overdetection. They characterize recall by assessing the accuracy of the predicted instances that are closest to the ground truth, but do not assess precision by penalizing the model for predicting additional instances with incorrect gaze targets. Additionally, the matching algorithm does not enforce that the ground truth instance is matched to a detection with a similar predicted head bounding box; if the heatmap loss dominates the matching cost, an instance may be selected based only on similarity between the predicted gaze heatmap and ground truth (see Fig 6 for examples). Thus, the model does not need to correctly associate people with their respective gaze targets to achieve high performance. For these reasons, the gaze metrics in this evaluation setting are not a direct comparison against our work and prior methods that follow the traditional problem formulation.
@@
We show the reported results of the 3 detection-based methods and our results on GazeFollow in Tab. 14. To quantitatively characterize the difference in evaluation setting, we also re-evaluate Tonini et al.’s [65] method on GazeFollow with the ground truth gaze label removed from the matching cost, using their published codebase. We alter the matching cost from Equation 6 to exclude the ground truth gaze label. This altered matching cost is defined as:
@@
We visualize the output instances of Tonini et al.’s default matching algorithm that uses ground truth gaze as part of the cost function, and our altered matching algorithm that does not use ground truth gaze in Fig. 6. The first two rows demonstrate cases where the matching algorithm chooses an instance with a predicted bounding box that is not associated with the correct person; the heatmap loss dominates the matching cost. With our altered matching function, an instance with a predicted bounding box for the correct person but incorrect gaze heatmap is retrieved. The third row shows an example of overdetection, where multiple instances are predicted with a head bounding box for the correct person, but different gaze targets. With the use of ground truth gaze during matching, the instance with the most correct heatmap is selected. However, without this ground truth information, the model does not select the best instance and produces an incorrect gaze prediction. These examples visually illustrate the difference in evaluation setting: when ground truth gaze information is used at test time, a model can achieve high performance by producing instances that capture different potential gaze targets and relying on the matching algorithm to retrieve the best instances to evaluate with. However, the gaze metrics do not characterize the model’s ability to determine which of these instances are indeed gaze targets and associate them with the correct person.